OpenStudy (anonymous):

The sum of two consecutive integers is between 5 and 9. What are the two integers? Tell the two numbers only

OpenStudy (watchmath):

2+3=5,3+4=7,4+5=9. So it must be 3 and 4

OpenStudy (anonymous):

Set up an inequality. 5 < x + x + 1< 9